OpenStudy (anonymous):

Find the surface area of a sphere with diameter 25 mm. A. 625 mm2 B. 157.08 mm2 C. 1963.50 mm2 D. 490.87 mm2

OpenStudy (anonymous):

B. 157.08 mm2

OpenStudy (anonymous):

0r 490.87 mm2

OpenStudy (callisto):

i guess it's the forth time i wrote this formula surface area = 4πr^2 and r = diameter /2 So can you work it out?

OpenStudy (experimentx):

4*pi*(25/2)^2 = 625*3.1514 = C

OpenStudy (king):

C.1963.5

OpenStudy (king):

@Rohangrr what are u doin u are giving only wrong answers check ure answers before u give them....